OpenStudy (anonymous):

solve using substitution? I have no idea how to do this ._. x=y-1 and -x+y=-1

OpenStudy (anonymous):

x=y-1 and -x+y=-1 Plug in x = y-1 into second equation -(y-1)+y=-1 y+1+y=-1 2y+1=-1 2y = -2 y=-1 x = -1 - 1 = -2 (-2, -1)

OpenStudy (anonymous):